OpenStudy (praxer):

What is the greatest power of 18 in 50C25

OpenStudy (praxer):

Mere ko dur dur tak koi idea nahin hain ki ye kiya hain.. please help :)

OpenStudy (praxer):

@ParthKohli

OpenStudy (praxer):

@satellite73

OpenStudy (praxer):

help me..... please !

OpenStudy (anonymous):

\[^{50} C_{25} = \frac{50 !}{25! \times 25!}\]

OpenStudy (anonymous):

By 50C25, this is what you mean??

OpenStudy (praxer):

I know this but I don't know WHAT THE POWER OF 18 IN 50C25 MEAN ?

OpenStudy (praxer):

@Luigi0210

OpenStudy (praxer):

@morgcarr12

OpenStudy (anonymous):

WHY ARE SO MANY PEOPLE TAGGING ME IM SO CONFUSED, lol ill try to figure it out tho

ganeshie8 (ganeshie8):

\(\large 18 = 2\times 3^2\) so the power of 3 is the limiting factor here

ganeshie8 (ganeshie8):

also notice that 50C25 is an integer because it represents number of ways of choosing 25 objects from 50 objects which is countable

ganeshie8 (ganeshie8):

Next, do you know how to get the highest exponent of 3 in the prime factorization of 50! ?

OpenStudy (praxer):

NO, :(

ganeshie8 (ganeshie8):

50! = 1.2.3.4.5.6.7.8...49.50

ganeshie8 (ganeshie8):

how many numbers less than 50 are multiples of 3 ?

OpenStudy (praxer):

15

OpenStudy (praxer):

OKAY ! WHAT TO DO FOR THE REST. :)

ganeshie8 (ganeshie8):

try again

ganeshie8 (ganeshie8):

number of multiples of 3 less than or equal to 50 = \(\large \left\lfloor\dfrac{50}{3}\right\rfloor = 16\)

OpenStudy (praxer):

3,6,9,12,15,18,21,24,27,30,33,36,39,42,45= 15 nos only.

OpenStudy (praxer):

oha ya 48 =1 silly me :(

ganeshie8 (ganeshie8):

what about 48 ?

ganeshie8 (ganeshie8):

:)

ganeshie8 (ganeshie8):

so can we say the exponent of 3 in prime factorization of 50! is 16 ?

ganeshie8 (ganeshie8):

nope, because multiples of 3^2's contribute two 3's, but we have counted them only once. so lets count how many multiples of 3^2's are there

ganeshie8 (ganeshie8):

how many multiples of 3^2's are there ?

OpenStudy (praxer):

so, the exponent of a number let x in y! is all the multiple of x less than y right? I don't know this, so inquired ?

ganeshie8 (ganeshie8):

consider a smaller example : 4! = 24 = 2^3*3^1

ganeshie8 (ganeshie8):

here the exponent of 2 is 3 and the exponent of 3 is 1

OpenStudy (praxer):

in 18 exponent of 3 is 2, than why did we find that for 50! Please explain me this.

ganeshie8 (ganeshie8):

we're trying to find out the exponent of 3 in 50!

ganeshie8 (ganeshie8):

because we want to see how many 3^2's are there in 50C25

OpenStudy (praxer):

okay let me conclude till here, we first expressed 18 in exponential of primes, we found it to be 3^2*2 now since we need to find the power of 18 in 50C25 so, we are finding the power of 3 in 50 and power of 3 in 25 correct.

ganeshie8 (ganeshie8):

our idea is this : 1) find the exponent of of 3 in 50! 2) find the exponent of of 3 in 25! play with above to figure out the exponent of 3^2 in 50C25

ganeshie8 (ganeshie8):

you're right!

ganeshie8 (ganeshie8):

lets finish finding the exponent of 3 in 50! first

OpenStudy (praxer):

okay so, for 50! we will find the multiples of 3^2 less than 50 or the multiples of 3 less than 50. Cause in one our last post you said this, I could not figure out what you said, "nope, because multiples of 3^2's contribute two 3's, but we have counted them only once. so lets count how many multiples of 3^2's are there"

OpenStudy (anonymous):

Number of 3s in 50! is: \[\frac{50}{3} + \frac{50}{3^2} + \frac{50}{3^3 } = 16 + 5 + 1\]

ganeshie8 (ganeshie8):

while counting multiples of 3, we did not account for numbers that have two 3's in them right ? for example, the numbers 9 and 18 are counted only once however they each contribute two 3's

OpenStudy (praxer):

okay so according to #waterineyes we have 22 3 in 50. than ?

ganeshie8 (ganeshie8):

similarly count number of 3's in 25!

OpenStudy (praxer):

25/3+25/3^2+25/3^3 this way ?

ganeshie8 (ganeshie8):

exactly !

ganeshie8 (ganeshie8):

take the integer value

OpenStudy (praxer):

measn if the value is 24.07 than i should take 24 ?

OpenStudy (praxer):

#means

ganeshie8 (ganeshie8):

yes

OpenStudy (praxer):

then ?

ganeshie8 (ganeshie8):

how many 3's are there in 25! ?

OpenStudy (praxer):

12.03 so, it will be 12. for 50 it is 24 not 22

ganeshie8 (ganeshie8):

nope

ganeshie8 (ganeshie8):

you need to take integer value for each fraction

OpenStudy (praxer):

ya for 50! according to the formula I found the value to be 24.07 than ?

OpenStudy (anonymous):

\[\frac{25}{3} + \frac{25}{3^2} = 8 + 2 = 10\]

ganeshie8 (ganeshie8):

exponent of 3 in 50! : \[\large \left\lfloor\dfrac{50}{3}\right\rfloor +\left\lfloor\dfrac{50}{3^2}\right\rfloor +\left\lfloor\dfrac{50}{3^3}\right\rfloor = 16+5+1 = 22\]

OpenStudy (praxer):

oho for each fraction i didnt notice that again.. Dhet teriki.. Silly me :(

OpenStudy (praxer):

okay so the value is 22+10=32

OpenStudy (anonymous):

It will be simple for you to understand if you take it in another way..

OpenStudy (anonymous):

Firstly divide 50 by 3, and tell me what did you get??

ganeshie8 (ganeshie8):

exponent of 3 in 25! : \[\large \left\lfloor\dfrac{25}{3}\right\rfloor +\left\lfloor\dfrac{25}{3^2}\right\rfloor +\left\lfloor\dfrac{25}{3^3}\right\rfloor = 8+2+0 = 10\]

OpenStudy (praxer):

okay i got it so, now for 50 it is 22 and for 25 it is 10 then ?

OpenStudy (anonymous):

Do not take the remainder, leave the remainder.. take only the quotient..

OpenStudy (praxer):

Okay got till here..

ganeshie8 (ganeshie8):

yes next lets see how can we use the information that we have to figure out the exponent of 3 in 50C25

ganeshie8 (ganeshie8):

whats 50C25 again ?

OpenStudy (praxer):

50!/25!*25!

OpenStudy (anonymous):

When you divide 50 by 3, you will get 16 as quotient.. Now, divide this 16 again by 3, you will get 5.. Now divide 5 by 3, you will get 1.. Now you cannot divide now 1 by 3, so you get now : 16 + 5 + 1 = 22

OpenStudy (praxer):

okay that seems good @waterineyes

ganeshie8 (ganeshie8):

can we write that like this : \[\large ^{50} C_{25} = \frac{50 !}{25!* 25!} = \dfrac{3^{22} a}{3^{10}b*3^{10}b}\] ?

OpenStudy (praxer):

so we have 3^2*a/b^2 understood till here. than ? :)

ganeshie8 (ganeshie8):

then we're done ! there is only one 9 in 50C25 what does that tell us ?

OpenStudy (anonymous):

Similarly you can do it for 25.. 25/3 = 8 8/3 = 2 As 2 < 3, so, now add : 8 + 2 = 10

OpenStudy (praxer):

it tells us that the answer is not matching with the solution where it is given to be 1

ganeshie8 (ganeshie8):

:/

OpenStudy (praxer):

:(

ganeshie8 (ganeshie8):

think a bit, there is a one 9 as factor in 50C25 : 50C25 = 9*something

OpenStudy (praxer):

ya 9*a/b^2

ganeshie8 (ganeshie8):

that precicely means the highest exponent of 18 is 1, right ?

OpenStudy (praxer):

how come the last statement is depicted.

ganeshie8 (ganeshie8):

because there will always be more number of 2's than 3's

ganeshie8 (ganeshie8):

since only one 9 is available, you can pull out only one 18 out of 50C25

ganeshie8 (ganeshie8):

if you want, you can find out the exponent of 2 as well..

OpenStudy (praxer):

can you provide me an algorithm to deal with such questions, seriously I have learnt it first time :) please. I mean how come the 18 at the last. :(

ganeshie8 (ganeshie8):

do you agree that the highest exponent of 3 in 50C25 is 2 ?

ganeshie8 (ganeshie8):

if so, next think about how many 3's are required for 18

OpenStudy (praxer):

since 50c25=9a/b^2 so the highest exponent shall be 2 only right ?

OpenStudy (praxer):

only thing I am stuck with is that, we were asked to find the power of 18 in 50C25 we fist prime factorised 18 and than we found the power of 3 in 50 and 25 and which we found to be 9a/b^2 than after this we can conclude that 1. highest power of 3 is 2 2. highest power of 18 is 2-------------- How ????????

OpenStudy (praxer):

sorry typo of 18 is 1 how ?

ganeshie8 (ganeshie8):

how many 3's are needed to cookup one 18 ?

ganeshie8 (ganeshie8):

and how many 3's are available in 50C25 ?

OpenStudy (praxer):

2 3's are needed.

OpenStudy (praxer):

3^2*2

ganeshie8 (ganeshie8):

yes

OpenStudy (praxer):

and we have 3^2*a/b^2 which means we have 2 3's . So we can get only 1 18, this is what you mean.

ganeshie8 (ganeshie8):

Exactly !

OpenStudy (praxer):

to aisa bolo na.. okay, so ever again if I am asked to find the power of some x in yCr than I shall do this 1. prime factorise x let it be 3*2 2. find the exponent of either 2 or 3 in 50 and than in 25 3. express the yCr term 4. find the number of 3 or 2 I have and match with number of 2 and 3 I require to make x yuhoo thats the answer number of x I can make. Correct ?